OpenStudy (anonymous):

I have to do improper integral comparison. What is larger than e^x / sin (x/2)?

OpenStudy (anonymous):

Your answer makes little sense, but remember sin is bounded.

OpenStudy (anonymous):

I apologise for not explaining it properly. My English has gone a little bit rusty, but your answer might actually prove to be useful. I'll use your advice and I'll see how it goes..

OpenStudy (anonymous):

\[\int\limits_{0}^{1} (e^{x})/(\sin(x/2))\]

OpenStudy (anonymous):

\[=\lim_{t\rightarrow0^+}\int_t^1{\frac{e^x}{\sin(\frac{x}{2})}}\]

OpenStudy (anonymous):

To use the comparison theorem, select an easier function like \[\frac{1}{\sin(x)}\]

OpenStudy (anonymous):

Thank you, agdgdgdgwngo! :)

OpenStudy (anonymous):

since that easier function is divergent on the interval (0, 1], then the original function is also divergent